OpenStudy (anonymous):

given the linear equation y=5x -1, find the y- coordinates of the points(-1), (2), and (1). Plot these points and graph linear equation.

OpenStudy (anonymous):

Simply plug those values in where x is and obtain 3 ordered pairs.

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

Something like this: y = 5(-1) - 1 y = 5(2) - 1 y = 5(1) - 1 Solve each of these.

OpenStudy (anonymous):

oh i got it thanks

OpenStudy (anonymous):

No problem, the results of those equations will be the y-value for each corresponding x. Graph the points and draw the line through them.

OpenStudy (anonymous):

thank you